write the general equation form of the line which has slope 3/4 and passes
through the point(10,-2)
I am having some trouble beginning this problem...i am not sure if a chart is needed in order to graph this...please help...thank you

Answer by foleysdad(7)

A linear equation has the form of y=mx+b where m = the slope of the line and b = the y intercept. Right now we know the slope 3/4 and we know one point that the line passes through (10,-2). Since the slope is defined as the rise over the run also known as the change in y over the change in x. This can be stated as (y1-y2)/(x1-x2) for the points (x1,y1) and (x2,y2). In our case our points are (10,-2) and (x2,y2) so the equation would look like
(-2-y2)/(10-x2)=3/4 since we are looking for the y intercept we know the x coordinate would zero, so the equation now looks like
(-2-y2)/(10-0)=3/4 Now lets simplify by multiplying both side by 10
(-2-y2)=30/4 Now multiply both sides by 4
4(-2-y2)=30 Simplify
-8-4y2 =30 Add 8 to both sides
-4y2=38 Divide both sides by -4
y2=-38/4 or y2=-9.5 Therefore the equation would be as follows
y = (3/4)x - 9.5